# LLM-PBE: Assessing Data Privacy in Large Language Models

### v1

@@ 1. Introduction
Despite these concerns, there exists a notable gap in the current research landscape: a lack of systematic analysis regarding the privacy of LLMs. Existing studies (Perez and Ribeiro, 2022; Mireshghallah et al., 2023; Zhang and Ippolito, 2023; Wei et al., 2023; Wang et al., 2023a; Pan et al., 2020) have the following limitations: 1) Limited evaluated data types: While the deployment of LLMs involves multiple stages and different types of data, most studies (Wang et al., 2023a; Wei et al., 2023; Zhang and Ippolito, 2023) only consider the potential leakage of a single type of data (e.g., Personally identifiable information (PII), prompts); 2) Limited models: While there are a rich set of LLMs currently, many analyses (Perez and Ribeiro, 2022; Zhang and Ippolito, 2023; Wei et al., 2023) are constrained to a few LLMs or smaller models such as GPT-2. 3) Limited attack approaches: Existing studies usually only consider a single attack method (e.g., data extraction attack (Carlini et al., 2019; Mireshghallah et al., 2022c)) and do not cover a broad range of attack metrics; 4) Limited consideration of privacy protection approaches: Existing studies (Perez and Ribeiro, 2022; Mireshghallah et al., 2023; Zhang and Ippolito, 2023; Wei et al., 2023; Wang et al., 2023a; Pan et al., 2020) usually lack the consideration of the effect of using privacy protection approaches on the data leakage. In summary, while these studies have touched upon specific aspects of privacy risks, a comprehensive evaluation encompassing the diverse facets of LLMs’ data privacy implications remains largely unexplored. This gap is evident in the fragmented approach of existing research, which often fails to consider the multi-dimensional nature of privacy risks in LLMs.
@@ 2.2. Data Privacy Leakage in LLMs
Data privacy in the context of LLMs concerns the protection of sensitive information that these models might access, learn, and potentially disclose during their operation. This encompasses personal data, confidential information, and any content that, if exposed, could lead to privacy breaches. The challenge in ensuring data privacy in LLMs arises from their training process, which involves large-scale datasets that can contain such sensitive information. Ensuring that these models respect user privacy and adhere to data protection standards is thus a critical concern. While developers usually provide inference services to LLMs without detailed information on the data collection and processing, numerous studies (Carlini et al., 2022b, 2021; Yu et al., 2023b; Panda et al., 2023b; Jayaraman et al., 2022) have shown that sensitive data may leak by just prompting LLMs as demonstrated in Figure 1. Thus, it is important to systematically assess the data privacy risks of LLMs.
@@ 2.3. Privacy Assessment of LLMs
As detailed in Table 1, current research in the field typically evaluates the privacy of LLMs using a limited range of models, datasets, and attack methodologies. For example, DecodingTrust (Wang et al., 2023a) evaluates the trustworthiness in GPT models on many aspects such as robustness, fairness, and privacy. However, for the privacy part, it only evaluates GPT models with a single attack method using different prompting context lengths. It finds that GPT-4 leaks more data than GPT-3.5, while our study aims to systematically compare different series of LLMs (e.g., Llama and GPTs) with different factors. Pan et al. (2020) demonstrate the privacy risks of language models assuming that the adversary has access to the text embedding, which does not fit in the current era of LLMs as adversaries usually do not have access to the embedding of training data. There are also many studies (Perez and Ribeiro, 2022; Mireshghallah et al., 2023; Zhang and Ippolito, 2023; Wei et al., 2023) that attack LLMs to demonstrate the existence of data leakage, but they focus on proposing a single attack/defend method instead of systematically benchmarking the privacy of LLMs to reveal the insights related to data privacy.
@@ 2.4. Privacy Enhancing Technologies for LLMs
There have been many data privacy protection approaches (Xiao and Tao, 2006; Bayardo and Agrawal, 2005; Xiao et al., 2010; Agrawal and Srikant, 2000). One popular approach is differential privacy (DP) (Dwork, 2006; Xiao et al., 2010; Xu et al., 2013; Dwork et al., 2014), which guarantees that the output does not change with a high probability even though an input data record changes. DP has been used in the training of machine learning models (Abadi et al., 2016; Phan et al., 2017; Shokri and Shmatikov, 2015), which is usually achieved by adding noises to gradients when using stochastic gradient descent. While using DP to retrain LLMs requires massive computing resources, it is possible to use DP to fine-tune LLMs as we will demonstrate in Section 3.6.2 and Section 4.4. Besides DP, we also exploit the potential usage of scrubbing (Pilán et al., 2022), machine unlearning (Jang et al., 2022; Wang et al., 2023b; Warnecke et al., 2021), and defensive prompting (yc-, 2023) for the data privacy protection in LLMs, which we will introduce in Section 3.6.
@@ 3.5.1. Data Extraction Attacks
We conclude that there are mainly two kinds of DEAs: query-based methods (inference-time attack) (Carlini et al., 2019, 2021; Mireshghallah et al., 2022c) and poisoning-based methods (training-time attack) (Jayaraman et al., 2022; Panda et al., 2023b). Query-based DEAs typically query LLMs to make them output training data. Poisoning-based methods modify the training data to insert poisons with a similar pattern as the target secret, and then easily extract this secret during inference. Since poisoning-based DEAs have a strong assumption that the attacker can access the training data, we only consider the query-based method in our toolkit. Specifically, we adopt the query-based method that prompts model with training data prefixes (Carlini et al., 2023a) (e.g., query ‘to: Alice ¡’ to make LLMs output the email address of Alice), and further explore different decoding configurations following (Yu et al., 2023b).
@@ 3.5.4. Jailbreaking Attacks
Like PLAs, JA prompts also have manually designed prompts and model-generated prompts. For manually designed prompts, we incorporate 15 JA prompting templates from public resources such as websites and papers  (Jai, 2023; Wei et al., 2023; Kang et al., 2023; Li et al., 2023a), which bypass the embedded safety requirements by obfuscating the input prompts or restricting the output format. For model-generated prompts, we use an existing approach (Chao et al., 2023) to generate the JA prompts using LLMs. Specifically, it uses one LLM to generate prompts, while using another LLM to judge whether the generated prompt successfully jailbreaks the target model. The generated prompts and responses are appended to the attack prompts in each round until successful jailbreaking.
@@ 3.6.2. Differential Privacy
DP has been applied in the training of machine learning models to protect training data (Abadi et al., 2016). However, since the training of LLMs requires a long time with massive computing resources, it is not feasible for us to use DP to retrain an LLM. Thus, we consider the usage of DP with parameter-efficient fine-tuning approaches such as LoRA (Hu et al., 2021). Instead of fine-tuning the whole model, we use LoRA to only fine-tune additional parameters with DP, whose size is much smaller than the size of LLM.
@@ 3.6.3. Machine Unlearning
While LLMs memorize some private training data, a promising way to protect data privacy is to update the model to unlearn specific data, i.e., machine unlearning. Machine unlearning has been an attractive research direction recently as data regulations such as GDPR stipulate that individuals have the “right to be forgotten”. While many machine learning studies are for computer vision (Tarun et al., 2023; Lin et al., 2023; Zhang et al., 2022), machine unlearning approaches for LLMs remain underexploited. Some studies (Jang et al., 2022; Wang et al., 2023b; Warnecke et al., 2021) fine-tune the trained model to unlearn the deleted data, which is more practical than modifying the training process (Bourtoule et al., 2021; Kumar et al., 2022) as the training of LLMs is very expensive. In our toolkit, we adopt an approach (Wang et al., 2023b) to fine-tune the LLM using knowledge gap alignment. Specifically, the LLM is updated such that the knowledge gap between it and the model trained on the deleted data is similar to the gap of another model handling the seen and unseen data.
@@ 3.7. Efficiency
Efficiency is an important factor that influences the practicality and scalability of various attack and defense strategies. Details regarding the GPU memory requirements and computational costs of these strategies are presented in Table 2. These experiments were conducted using the Llama-2 7B model on the Enron dataset, utilizing a system equipped with two NVIDIA A100 GPUs and two AMD EPYC 7J13 64-Core Processors. The findings indicate that most attack strategies are computationally efficient as they do not necessitate the training or updating of models. However, model-based MIAs are not feasible for LLMs due to the necessity of training multiple LLMs to develop an effective attack model. Among the defense mechanisms, DP-SGD offers lower computational overhead compared to data scrubbing. This is because DP-SGD integrates minimal additional operations into the training process, whereas scrubbing requires extensive preprocessing of the original data using language models. Despite all approaches needing at least 28GB of GPU memory owing to the large parameter sizes involved, the availability of LLM inference services from various companies (e.g., OpenAI, TogetherAI) means that attackers might not require local model hosting, potentially reducing the need for high-performance GPUs.
Table 2. The peak GPU memory (GB) and computational overhead per sample for the attack/defense methods.
GPU mem (GB) Cost
DEAs Query-based 33 27s
Poison-based 56 28s
MIAs Model-based
Comparison-based 33 2.5s
PLAs Manually-designed 30 2.1s
Model-generated 34 16m
JAs Manually-designed 29 1.8s
Model-generated 36 12m
PETs Scrubbing 11 2.1h
DP-SGD 112 26m
@@ 4. Leakage of Training Data
In this section, we conduct extensive experiments to assess the privacy of training data of LLMs with existing attack methods, including data used for pertaining and fine-tuning. We focus on answering the following research questions: 1) Does the privacy risks of in LLMs correspond proportionally with their increasing scale and effectiveness? 2) How are different data characteristics associated with the privacy risks of LLMs? 3) Are there practical privacy-preserving approaches when deploying LLMs? Due to the page limit, we present representative experiments in the main paper and put additional results in Appendix.
@@ 4.1. Experimental Setup
Datasets We evaluate the following datasets including 1) Enron (Klimt and Yang, 2004) dataset that contains 500k emails generated by employees of the Enron Corporation; 2) ECHR (Chalkidis et al., 2019) dataset that contains 11.5k cases from the European Court of Human Rights; 3) Github dataset where we collect the Python code from 22k repositories in Github that have stars over 500. For the results on Github, please refer to Appendix C.2.
@@ 4.2. Effect of Model Size
As demonstrated in existing studies (Staab et al., 2024; Treutlein et al., 2024), LLMs can also infer private information from the input context. To investigate whether memorization or reasoning primarily contributes to DEAs, we also conduct DEAs on a synthetic email dataset that the model has never seen (as shown in DEA Synthetic). From the results, we observe that DEA accuracy is zero in most cases, indicating that the model is not able to infer complete email addresses accurately through reasoning. Thus, LLMs indeed memorize training data, which poses potential privacy risks.
Takeaways: Within the same series of LLMs trained on identical data in the same order, as the size of the models increases, their capacities on language tasks also increase. Concurrently, these larger models exhibit enhanced extraction accuracy with existing DEAs, due to their advanced memorization capacities. Notably, the rate of increase in data extraction accuracy outpaces the improvements in model utility, suggesting a growing privacy risk as models scale.
@@ 4.3. Effect of Data Characteristics
Data type. To investigate the effect of data type on privacy risks, we use DEAs with ECHR dataset on Llama-2 7b (Touvron et al., 2023), which includes different types of PII types including name, location, and date. The results are shown in Figure 5. The proportions of samples of name, location, and date are 43.9%, 9.7%, and 46.4%, respectively. From the figure, it is evident that text data (i.e., name and position) is more susceptible to leakage than digit data (i.e., date). The contextual richness of text data in training sets facilitates easier learning and recall by the model. This rich context offers numerous ‘hooks’ for the model to engage with, unlike the more isolated and context-free nature of digit data, enhancing the model’s propensity to retain and subsequently leak textual information.
Figure 5. DEA accuracy of different positions and types of data on ECHR.
Position of Private Data. Following the above setup, we explore how the position of private within a sentence — whether at the beginning, in the middle, or at the end — impacts the accuracy of DEA. The results are presented in Figure 5. The proportions of samples in front, middle, and end are 25.1%, 36.5%, and 38.4%, respectively. We observe that private data that appears in the earlier position of a sentence has a higher data extraction accuracy. In transformer-based LLMs, the attention mechanism tends to focus more heavily on the important part of a sentence (Vaswani et al., 2017). When private data appears at the beginning, we suspect that it is more likely to be captured and emphasized by the model’s attention layers, making it more susceptible to extraction.
@@
Takeaways: Our findings reveal that data type, data position, data length, and pretraining data size collectively impact privacy risks on Llama-2. Textual data is more susceptible to leakage compared to numerical data due to its contextual richness. Private data at the beginning of a sentence is more vulnerable to extraction by the model’s attention mechanisms. Data samples that are harder to predict, indicated by higher perplexity, are more easily identified in MIAs. Additionally, increasing the size of the training data enhances the model’s memorization capacity, leading to higher privacy risks. These insights highlight the necessity for targeted privacy strategies that address the specific characteristics of different data types in LLMs.

### v2

@@ 1. Introduction
Despite these concerns, there exists a notable gap in the current research landscape: a lack of systematic analysis regarding the privacy of LLMs. Existing studies (Perez and Ribeiro, 2022; Mireshghallah et al., 2024; Zhang and Ippolito, 2023; Wei et al., 2023; Wang et al., 2023a; Pan et al., 2020) have the following limitations: 1) Limited evaluated data types: While the deployment of LLMs involves multiple stages and different types of data, most studies (Wang et al., 2023a; Wei et al., 2023; Zhang and Ippolito, 2023) only consider the potential leakage of a single type of data (e.g., Personally identifiable information (PII), prompts); 2) Limited models: While there are a rich set of LLMs currently, many analyses (Perez and Ribeiro, 2022; Zhang and Ippolito, 2023; Wei et al., 2023) are constrained to a few LLMs or smaller models such as GPT-2. 3) Limited attack approaches: Existing studies usually only consider a single attack method (e.g., data extraction attack (Carlini et al., 2019; Mireshghallah et al., 2022c)) and do not cover a broad range of attack metrics; 4) Limited consideration of privacy protection approaches: Existing studies (Perez and Ribeiro, 2022; Mireshghallah et al., 2024; Zhang and Ippolito, 2023; Wei et al., 2023; Wang et al., 2023a; Pan et al., 2020) usually lack the consideration of the effect of using privacy protection approaches on the data leakage. In summary, while these studies have touched upon specific aspects of privacy risks, a comprehensive evaluation encompassing the diverse facets of LLMs’ data privacy implications remains largely unexplored. This gap is evident in the fragmented approach of existing research, which often fails to consider the multi-dimensional nature of privacy risks in LLMs.
@@ 2.2. Data Privacy Leakage in LLMs
Data privacy in the context of LLMs concerns the protection of sensitive information that these models might access, learn, and potentially disclose during their operation. This encompasses personal data, confidential information, and any content that, if exposed, could lead to privacy breaches. The challenge in ensuring data privacy in LLMs arises from their training process, which involves large-scale datasets that can contain such sensitive information. Ensuring that these models respect user privacy and adhere to data protection standards is thus a critical concern. While developers usually provide inference services to LLMs without detailed information on the data collection and processing, numerous studies (Carlini et al., 2022b, 2021; Yu et al., 2023b; Panda et al., 2024; Jayaraman et al., 2022) have shown that sensitive data may leak by just prompting LLMs as demonstrated in Figure 1. Thus, it is important to systematically assess the data privacy risks of LLMs.
@@ 2.3. Privacy Assessment of LLMs
As detailed in Table 1, current research in the field typically evaluates the privacy of LLMs using a limited range of models, datasets, and attack methodologies. For example, DecodingTrust (Wang et al., 2023a) evaluates the trustworthiness in GPT models on many aspects such as robustness, fairness, and privacy. However, for the privacy part, it only evaluates GPT models with a single attack method using different prompting context lengths. It finds that GPT-4 leaks more data than GPT-3.5, while our study aims to systematically compare different series of LLMs (e.g., Llama and GPTs) with different factors. Melis et al. (2019) and Pan et al. (2020) demonstrate the privacy risks of recovering texts by the text embeddings, which does not fit in the current era of LLMs as adversaries usually do not have access to the embedding of training data. There are also many studies (Perez and Ribeiro, 2022; Mireshghallah et al., 2024; Zhang and Ippolito, 2023; Wei et al., 2023) that attack LLMs to demonstrate the existence of data leakage, but they focus on proposing a single attack/defend method instead of systematically benchmarking the privacy of LLMs to reveal the insights related to data privacy.
@@ 2.4. Privacy Enhancing Technologies for LLMs
There have been many data privacy protection approaches (Xiao and Tao, 2006; Bayardo and Agrawal, 2005; Xiao et al., 2010; Agrawal and Srikant, 2000). One popular approach is differential privacy (DP) (Dwork, 2006; Xiao et al., 2010; Xu et al., 2013; Dwork et al., 2014), which guarantees that the output does not change with a high probability even though an input data record changes. DP has been used in the training of machine learning models (Abadi et al., 2016; Phan et al., 2017; Shokri and Shmatikov, 2015), which is usually achieved by adding noises to gradients when using stochastic gradient descent. While using DP to retrain LLMs requires massive computing resources, it is possible to use DP to fine-tune LLMs as we will demonstrate in Section 3.6.2 and Section 4.4. Besides DP, we also exploit the potential usage of scrubbing (Pilán et al., 2022), machine unlearning (Jang et al., 2023; Wang et al., 2023b; Warnecke et al., 2021), and defensive prompting (yc-, 2023) for the data privacy protection in LLMs, which we will introduce in Section 3.6.
@@ 3.5.1. Data Extraction Attacks
We conclude that there are mainly two kinds of DEAs: query-based methods (inference-time attack) (Carlini et al., 2019, 2021; Mireshghallah et al., 2022c) and poisoning-based methods (training-time attack) (Jayaraman et al., 2022; Panda et al., 2024). Query-based DEAs typically query LLMs to make them output training data. Poisoning-based methods modify the training data to insert poisons with a similar pattern as the target secret, and then easily extract this secret during inference. Since poisoning-based DEAs have a strong assumption that the attacker can access the training data, we only consider the query-based method in our toolkit. Specifically, we adopt the query-based method that prompts model with training data prefixes (Carlini et al., 2023a) (e.g., query ‘to: Alice ¡’ to make LLMs output the email address of Alice), and further explore different decoding configurations following (Yu et al., 2023b).
@@ 3.5.4. Jailbreaking Attacks
Like PLAs, JA prompts also have manually designed prompts and model-generated prompts. For manually designed prompts, we incorporate 15 JA prompting templates from public resources such as websites and papers  (Jai, 2023; Wei et al., 2023; Kang et al., 2024; Li et al., 2023a), which bypass the embedded safety requirements by obfuscating the input prompts or restricting the output format. For model-generated prompts, we use an existing approach (Chao et al., 2023) to generate the JA prompts using LLMs. Specifically, it uses one LLM to generate prompts, while using another LLM to judge whether the generated prompt successfully jailbreaks the target model. The generated prompts and responses are appended to the attack prompts in each round until successful jailbreaking.
@@ 3.6.2. Differential Privacy
DP Stochastic Gradient Descent (DP-SGD) has been widely applied in the training of machine learning models to protect training data (Abadi et al., 2016). As training LLMs requires a long time with massive computing resources (e.g., GPU memory) and DP-SGD further augment it with sample-wise gradient clipping, we leverage advanced memory-efficient DP-SGD techniques including automatic clipping and adaptive multiple-GPU distribute computation (Bu et al., 2023a, c, b, e) to reduce the resource load.
@@ 3.6.3. Machine Unlearning
While LLMs memorize some private training data, a promising way to protect data privacy is to update the model to unlearn specific data, i.e., machine unlearning. Machine unlearning has been an attractive research direction recently as data regulations such as GDPR stipulate that individuals have the “right to be forgotten”. While many machine learning studies are for computer vision (Tarun et al., 2023; Lin et al., 2023; Zhang et al., 2022), machine unlearning approaches for LLMs remain underexploited. Some studies (Jang et al., 2023; Wang et al., 2023b; Warnecke et al., 2021) fine-tune the trained model to unlearn the deleted data, which is more practical than modifying the training process (Bourtoule et al., 2021; Kumar et al., 2023) as the training of LLMs is very expensive. In our toolkit, we adopt an approach (Wang et al., 2023b) to fine-tune the LLM using knowledge gap alignment. Specifically, the LLM is updated such that the knowledge gap between it and the model trained on the deleted data is similar to the gap of another model handling the seen and unseen data.
@@ 3.7. Efficiency
Efficiency is an important factor that influences the practicality and scalability of various attack and defense strategies. Details regarding the GPU memory requirements and computational costs of these strategies are presented in Table 2. For comparison-based MIAs, we report the range of costs across PPL, LiRA (Mattern et al., 2023), and Refer (Carlini et al., 2021) (details about these approaches are introduced in Section 4.1). For other categories, we report the usual GPU memory requirements and costs. These experiments were conducted using the Llama-2 7B model on the Enron dataset, utilizing a system equipped with two NVIDIA H100 GPUs and four AMD EPYC 9654 96-Core Processors. The attack and defense approaches can be broadly categorized into two types: inference-based methods and training-based methods. Inference-based methods typically involve querying the models, making them efficient. In contrast, training-based methods usually require fine-tuning or training LLMs, which is significantly more costly. Model-based MIAs require the training of multiple models to mimic the target model, which is not feasible for LLMs. Despite all approaches requiring at least 28GB of GPU memory due to the large parameter sizes involved, the availability of LLM inference services from various companies (e.g., OpenAI, TogetherAI) means that attackers might not need to host models locally.
Table 2. The required GPU memory size (GB) and computational cost per sample for the attack/defense methods compared with the plain LLM inference and training.
GPU mem Cost
Inference-based methods Plain Inference 28GB 0.43s
Query-based DEAs 1x $\sim$1x
Comparison-based MIAs 1x 1-2x
PLAs 1x 1x
JAs 1x 1x
Training-based methods Plain Training 112GB 0.59s
Poison-based DEAs 1x 1x
Model-based MIAs
DP-SGD $\sim$1x 1.2x
Others Scrubbing 13.5GB 0.04s
@@ 4. Leakage of Training Data
In this section, we conduct extensive experiments to assess the privacy of training data of LLMs with existing attack methods, including data used for pertaining and fine-tuning. We focus on answering the following research questions: 1) Does the privacy risks of in LLMs correspond proportionally with their increasing scale and effectiveness? 2) How are different data characteristics associated with the privacy risks of LLMs? 3) Are there practical privacy-preserving approaches when deploying LLMs? We present representative experiments in the main paper and put additional results in Appendix.
@@ 4.1. Experimental Setup
Datasets We evaluate 1) Enron (Klimt and Yang, 2004) dataset that contains 500k emails generated by employees of the Enron Corporation; 2) ECHR (Chalkidis et al., 2019) dataset that contains 11.5k cases from the European Court of Human Rights; 3) Github dataset where we collect the Python code from 22k repositories in Github that have stars over 500. For the results on Github, please refer to Appendix C.2.
@@ 4.2. Effect of Model Size
As demonstrated in existing studies (Staab et al., 2024; Treutlein et al., 2024), LLMs can also infer private information from the input context. To investigate whether memorization or reasoning primarily contributes to DEAs, we also conduct DEAs on a synthetic email dataset that the model has never seen (as shown in DEA Synthetic), which has the same format as Enron. From the results, we observe that DEA accuracy is zero in most cases, indicating that the model is not able to infer complete email addresses accurately through reasoning. Thus, LLMs indeed memorize training data, which poses potential privacy risks.
Takeaways: Within the same series of LLMs trained on identical data in the same order, as the size of the models increases, their capacities on language tasks also increase. Concurrently, these larger models exhibit enhanced extraction accuracy with existing DEAs, due to their advanced memorization capacities. Notably, the rate of increase in data extraction accuracy on Enron outpaces the improvements in ARC-Easy for Pythia, suggesting a growing privacy risk as models scale.
@@ 4.3. Effect of Data Characteristics
Data type. To investigate the effect of data type on privacy risks, we use DEAs with ECHR dataset on Llama-2 7b (Touvron et al., 2023), which includes different types of PII types including name, location, and date. To ensure a fair comparison, we filter the ECHR dataset to include prompts with PIIs that only appear once. This filtered version is then used for DEAs. The proportions of samples of name, location, and date are 25.8%, 4.0%, and 70.2%. The results are shown in Table 3. While locations may be inferred by the LLM from the input context, the DEA for names and dates is very low on the plain Llama-2 7B model. Comparing the results between Llama-2 7B-FT and Llama2-7B, locations tend to be more susceptible to memorization during fine-tuning than dates. In the ECHR dataset, locations are more easily related to contextual information (e.g., the context includes the country of the location), which aids in the memorization of this information.
Table 3. DEA accuracy of different positions and types of data on ECHR. Llama-2 7B-FT is the Llama-2 7B model fine-tuned on ECHR with four epochs.
Model Type DEA (%) by position
Overall Front Middle End
Llama-2 7B name 0.81% 0.87% 0.58% 1.0%
location 2.6% 3.8% 2.5% 2.3%
date 0.30% 0.34% 0.28% 0.30%
Llama-2 7B-FT name 10.4% 4.3% 12.7% 10.8%
location 19.2% 7.7% 17.3% 24.4%
date 6.7% 3.2% 5.3% 9.7%
Position of Private Data. We also explore how the position of private within a sample — whether at the beginning, in the middle, or at the end — impacts the accuracy of DEA as shown in Table 3. The proportions of samples in front, middle, and end are 17.7%, 42.7%, and 39.6%, respectively. We observe that private data that appears at the end of a sample usually has a higher data extraction accuracy. In transformer-based LLMs, the attention mechanism tends to focus more heavily on the important part of a sample (Vaswani et al., 2017). When private data appears at the end, we suspect that it is more likely to be captured and emphasized by the model’s attention layers, making it more susceptible to extraction.
@@
Takeaways: Our findings reveal that data type, data position, data length, and pretraining data size collectively impact privacy risks on Llama-2. Data with richer contextual information (e.g., locations) tends to be more susceptible to memorization during fine-tuning. Private data at the end of a sample is more vulnerable to extraction. Data samples that are harder to predict, indicated by higher perplexity, are more easily identified in MIAs. Additionally, increasing the size of the training data enhances the model’s memorization capacity, leading to higher privacy risks. These insights highlight the necessity for targeted privacy strategies that address the specific characteristics of different data types in LLMs.